# Fast and Private Inference of Deep Neural Networks by Co-designing Activation Functions

### v1

@@ Abstract
Machine Learning as a Service (MLaaS) is an increasingly popular design where a company with abundant computing resources trains a deep neural network and offers query access for tasks like image classification. The challenge with this design is that MLaaS requires the client to reveal their potentially sensitive queries to the company hosting the model. Multi-party computation (MPC) protects the client’s data by allowing encrypted inferences. However, current approaches suffer prohibitively large inference times. The inference time bottleneck in MPC is the evaluation of non-linear layers such as ReLU activation functions. Motivated by the success of previous work co-designing machine learning and MPC aspects, we develop an activation function co-design. We replace all ReLUs with a polynomial approximation and evaluate them with single-round MPC protocols, which give state-of-the-art inference times in wide-area networks. Furthermore, to address the accuracy issues previously encountered with polynomial activations, we propose a novel training algorithm that gives accuracy competitive with plaintext models. Our evaluation shows between $4$ and 90×90$\times$90 × speedups in inference time on large models with up to $23$ million parameters while maintaining competitive inference accuracy.
@@ 1 Introduction
[scale=0.9] {axis}[ title=, xlabel=Inference Time (s), ylabel=Test Accuracy (%), xmode=log, grid=major, ymin=88, ymax=96, legend pos=south east, ]
[…]
@@
This work focuses on reducing the runtime of secure inference on image data, under realistic network conditions, while maintaining classification accuracy. We consider the two-party setting using multi-party computation (MPC), where the server holds the ML model, and the client holds the data to query the model. Recent state-of-the-art works in this space employ various co-design approaches to reduce the inference time [30]. For example, COINN co-designs ML models optimized for quantization with efficient MPC protocols tailored to the custom models [17]. COINN substantially compresses the model and makes numerous optimizations to the architecture to achieve fast inferences. Another example is GForce, which tailors the cryptography needed for ML to high-speed GPU hardware [29]. By offloading vast amounts of work to the pre-computation phase, they are able to achieve state-of-the-art runtime and accuracy in secure inference [29]. Similarly, CryptGPU [32] modifies the CrypTen framework [20] to run efficiently on the GPU and give state-of-the-art inference times in wide area networks. However, despite making major steps towards practical inference, none of these works remove a crucial bottleneck in secure inference: the non-linear layers.
@@
This work addresses the non-linear layers by taking a co-design approach between the activation functions and MPC. We take the approach of replacing classic ReLU activation functions with a polynomial approximation to avoid conversions altogether. Previous work has considered this approach but with limited success [12]. We propose two modifications to make this approach practical. First, we develop and evaluate new single-round MPC protocols that give the fastest evaluation of polynomials to date. The challenge with using polynomials is they severely impact model accuracy [17, 27, 12]. Previous work could not successfully train DNNs with more than $11$ layers due to exploding gradients [12]. Thus, our second contribution is tailoring the ML training process to ensure high accuracy and stable training using polynomials. Our approach utilizes a new type of regularization that focuses on keeping the input to each activation function within a small range. We achieve close to plaintext accuracy on models as deep as ResNet-110 [16] and as large as a ResNet-50 on ImageNet [10] (23 million parameters). The combination of these approaches yields a co-design with state-of-the-art inference times and the highest accuracy for polynomial models.
We compare our work with three solutions representing the state-of-the-art approaches in secure inference according to Ng and Chow [30]. We summarize our results in Figure 1. Combining the single-round MPC protocols with our activation regularization achieves significantly faster inference times than all other solutions. Specifically, our solution is faster than CryptGPU by 4×4$\times$4 ×, GForce by 5×5$\times$5 ×, and COINN by 18×18$\times$18 × on average in wide area networks. Our approach also scales to large models on ImageNet with a 90×90$\times$90 × speedup over COINN and a 3×3$\times$3 × speedup over CryptGPU. Furthermore, our inference accuracy remains competitive with all other solutions. CryptGPU often gives slightly higher accuracy as it can evaluate any plaintext model (albeit slower than our work). Thus, the challenge for future work is to further close the ML accuracy gap between plain and polynomial models.
@@ 2.2 Neural Network Inference
We consider DNN classifiers with domain $\mathcal{X}\subseteq\mathbb{R}^{d}$ and range $\mathcal{Y}\subseteq\mathbb{R}^{c}$. DNN classifiers consist of a sequence of layers, each performing either a linear or a non-linear operation. The ResNet [16] architecture we consider is composed of (i) convolutional, (ii) fully connected, (iii) pooling, (iv) batch normalization, and (v) ReLU layers. All layers are linear, except for $\text{ReLU}(x)=\text{max}(x,0)$ and max pooling (that can be replaced with average pooling). To classify an input $x$, a classifier $h$, passes the input sequentially through each layer. Upon reaching the last layer, the prediction is obtained by taking $\text{arg max}_{i\in\{1..c\}}h(x)_{i}$, where we call $h$ the logit function for a classifier $h:\mathcal{X}\rightarrow\mathcal{Y}$. The output of the secure inference protocol for an encrypted input $x$ is the encrypted output $h(x)$ of the logit function.
@@ 3.1 Problem Setup
We follow the same threat model as prior work for two-party secure inference [17, 20, 27, 29]. Specifically, we follow the two-party client-server model where the server has a machine learning model (a DNN) they have trained, and the client holds private data upon which they would like to make an inference. The server’s input to the protocol is the weights of their trained model, which they do not want to leak to the client (due to intellectual property or protecting their MLaaS business [33]). The client has a private input (typically an image) they would like to classify using the model but do not want to leak this input or the prediction to the server. That is, the MPC function can be written as $f(\text{image},\text{model})=(label,\emptyset)$. Following previous work [17, 20, 27, 29], we consider the semi-honest model, introduced by Goldreich [15, §7.2.2], where adversaries do not deviate from the protocol but may gather information to infer private information. Also, in line with previous work, we assume the model architecture is known to both parties. This includes the dimensions and type of each layer and parameters such as field size used for inference. The mean and standard deviation of the training set are also known to both parties following CrypTen [20].
@@ 3.2 Privacy During Model Training
We focus only on the inference phase of machine learning. However, the privacy of the training process and training data is an orthogonal but essential problem. We recommend that the data owner take appropriate steps to protect the privacy of the model, such as training using differential privacy [1] or rounding the output of the inference. Furthermore, during training, care should be taken to protect against threats such as model stealing, which can be launched using only the inference result [18]. To summarize, the model owner learns nothing other than the fact a query was made. We ensure only the inference is revealed to the client; however, ML attacks that only require black box query access [18] must be defended against during the training process.
@@ 3.3 Motivating the Co-Design of Activation Functions
Motivated by this bottleneck, several works have focused on either reducing the number of ReLUs or replacing ReLUs altogether [13, 12, 19, 22, 27, 28]. One approach is to approximate each ReLU with a high degree polynomial [12, 22]. The advantage of using polynomials is that polynomials can be computed in arithmetic shares, thus removing the need for expensive conversions and improving the total inference time. A significant challenge with polynomials is maintaining model accuracy [12, 17, 13, 19, 27].
@@ 4.2 ESPN: Exponentiating Secret Shared Values using Pascal’s triangle
We observe that, for each $i$ in the sum, party $a$ can compute $a_{i}=x_{a}^{k-i}$ without needing to communicate with party $b$ (Alg. 1 line 4). Similarly, party $b$ can compute $x_{b}^{i}$ without communicating with party $a$ (Alg. 1 line 5). Finally, ($k\choose i$ can be computed by any party (or pre-computed ahead of time). For simplicity, we assign the computation of ($k\choose i$ to party $b$. Thus party $b$ computes $b_{i}={k\choose i}x_{b}^{i}$.
Once each party has computed their respective vectors, we multiply $a_{i}\cdot b_{i}$ for each $i$ in parallel (Alg. 1 line 6). We carry out this multiplication using standard MPC protocols in one round. To use these multiplication protocols, each party must have a share of the input. We use a trivial additive secret sharing, where the other party inputs zero as their share to the protocol (Alg. 1 line 2). Finally, after the multiplication, the sum of the binomial theorem can be efficiently computed with no communication (Alg. 1 line 7).
@@
2:     $\mathbf{a}=[[\mathbf{0}^{k}]]$, $\mathbf{b}=[[\mathbf{0}^{k}]]$triangleright Initializing Shares
3:     for $i=1:k$ do
4:         Party $a$ computes: $a_{i}=a_{i}+x_{a}^{k-i}$
5:         Party $b$ computes: $b_{i}=b_{i}+{k\choose i}x_{b}^{i}$
6:     $\mathbf{p}=BeaverMultiply(\mathbf{a},\mathbf{b})$$\triangleright$ Parallel Multiplication
7:     $s=\sum_{i=1}^{k}\mathbf{p_{i}}$triangleright $s=\mathbf{a}\cdot\mathbf{b}$
@@
Using our exponentiation protocol, we can now efficiently compute high-degree polynomials in a single round. The first step in evaluating a polynomial is to compute all needed exponents of the input (i.e. $\{x^{k}|k\in[n]\}$ for a degree $n$ polynomial) by calling Algorithm 1 in parallel. We note there will be a significant overlap in each party’s local powers from the binomial theorem. Thus, a simple cache can improve performance significantly. Furthermore, the binomial coefficients (${k\choose i}$, can also benefit from basic dynamic programming by reusing the previous result ($k-1$) to compute the next ($k$). After computing the exponents, the output of the polynomial can be computed by multiplying the coefficients (public values) and summing, all of which can be done locally.
@@ 4.3 Alternative Single Round Protocol: HoneyBadger
Using dynamic programming, the parties can then compute any power ($x^{k}r^{0}$) using only additions of previously computed terms and powers of $r$. To compute polynomials using this protocol, we need to multiply the coefficients and sum the terms, precisely as we did in ESPN.
The advantage of Lu et al.’s protocol is that the communication is small (only the opening of $x-r$). The first disadvantage is that the protocol requires a modified pre-computation phase, which is as difficult to pre-compute securely as the original problem (it is exponentiation). On the contrary, our binomial protocol uses standard beaver triplets commonly found in MPC frameworks. There are well established protocols for efficiently computing these triplets, and the parties may already have them due to the popularity of Beaver’s trick. The second disadvantage of Lu et al.’s solution is that, while the protocol requires very little communication, it is not locally parallelizable as each dynamic programming step depends on the previous one. In contrast, our entire protocol can be executed in parallel.
We also consider the runtime of using HoneyBadger to compute polynomial approximations of ReLUs in Figure 2. We emphasize this is a runtime-only evaluation. Without our training algorithm in Section 5, none of the polynomial-based solutions can attain usable accuracy. We find that ESPN and HoneyBadger perform similarly in practice, with HoneyBadger gaining a slight advantage in very low network delay. Due to the trade-offs and similarities, we will evaluate both approaches for the remainder of this work.
4.4 Floating Point Considerations
We note that the exponentiation protocols we have discussed are designed for integers. Extending to floating point values is straightforward, but requires rescaling (a standard practice in fixed-point arithmetic). Furthermore, in both protocols (ESPN and HoneyBadger), we can only rescale once for each polynomial computation (since the majority of the computation is local). Thus, to ensure correctness, we must scale up each exponent by $n+1-k$ where $n$ is the degree of the polynomial and $k$ is the exponent (Algorithm 1). This scaling up is a scalar multiplication by a public constant and thus incurs no additional rounds. At the end of the computation, we must then scale down by $n$ for correctness. We use CrypTen’s two-party truncation protocol that also incurs no additional rounds. However, there is a negligible chance of an incorrect result from this truncation protocol due to wrap-around in the ring. Specifically, the probability of an incorrect result when truncating $x$ is $\frac{x}{2^{L}}$ where $2^{L}$ is the size of the ring[20]. This implies $x$ must be small compared to the ring for this fast truncation protocol to be correct. We observe that $x$ can also be very large while maintaining correctness due to symmetry. Following the proof of Mohassel and Zhang [28], we get that the probability of failure is $max\{\frac{x}{2^{L}},\frac{2^{L}-x}{2^{L}}\}$. Since we upscale all terms by $n+1-k$ we ensure $x$ is sufficiently large and thus can take advantage of this fast truncation protocol with no additional rounds.
@@ 4.4 Floating Point Considerations
In addition to scaling, we must ensure that the polynomial computation does not overflow in the ring by choosing the correct precision. Consider using degree $n$ polynomials fitted to a range $[-\lambda,\lambda]$. If the global precision (size of the ring) is $L$-bit, then the working precision of each value must be $p$-bit where
$\log{\left\lceil n\cdot\frac{1-\lambda^{n}}{1-\lambda}\right\rceil}+n\cdot p\leq L-1$ (5)
For our experiments, we use CrypTen, with $L=64$ [20]. Assuming default values of $n=4$ and $\lambda=5$ we get that $10$-bit is the maximum working precision. However, this is a pessimistic upper bound, and in practice, we find that $12$-bit precision gives the best results.

### v2

@@ Abstract
Machine Learning as a Service (MLaaS) is an increasingly popular design where a company with abundant computing resources trains a deep neural network and offers query access for tasks like image classification. The challenge with this design is that MLaaS requires the client to reveal their potentially sensitive queries to the company hosting the model. Multi-party computation (MPC) protects the client’s data by allowing encrypted inferences. However, current approaches suffer from prohibitively large inference times. The inference time bottleneck in MPC is the evaluation of non-linear layers such as ReLU activation functions. Motivated by the success of previous work co-designing machine learning and MPC, we develop an activation function co-design. We replace all ReLUs with a polynomial approximation and evaluate them with single-round MPC protocols, which give state-of-the-art inference times in wide-area networks. Furthermore, to address the accuracy issues previously encountered with polynomial activations, we propose a novel training algorithm that gives accuracy competitive with plaintext models. Our evaluation shows between $3$ and 110×110$\times$110 × speedups in inference time on large models with up to $23$ million parameters while maintaining competitive inference accuracy.
@@ 1 Introduction
$10^{1}$$10^{2}$8888888890909090929292929494949496969696Inference Time (s)Test Accuracy (%)This WorkCOINN [18]GForce [31]CryptGPU [35]
@@
This work focuses on reducing the runtime of secure inference on image data, under realistic network conditions, while maintaining classification accuracy. We consider the two-party setting using multi-party computation (MPC), where the server holds the modified ML model, and the client holds the data to query the model. Recent state-of-the-art works in this space employ various co-design approaches to reduce the inference time [32]. For example, COINN co-designs ML models optimized for quantization with efficient MPC protocols tailored to the custom models [18]. COINN substantially compresses the model and makes numerous optimizations to the architecture to achieve fast inferences. Another example is GForce, which tailors the cryptography needed for ML to high-speed GPU hardware [31]. By offloading vast amounts of work to the pre-computation phase, they are able to achieve state-of-the-art runtime and accuracy in secure inference [31]. Similarly, CryptGPU [35] modifies the CrypTen framework [22] to run efficiently on the GPU and give state-of-the-art inference times in wide area networks. However, despite making major steps towards practical inference, none of these works remove a crucial bottleneck in secure inference: the non-linear layers.
@@
This work addresses the non-linear layers by taking a co-design approach between the activation functions and MPC. We take the approach of replacing classic ReLU activation functions with a polynomial approximation during training and inference to avoid conversions altogether. Previous work has considered this approach but with limited success [12]. We propose two modifications to make this approach practical. First, we develop and evaluate new single-round MPC protocols that give the fastest evaluation of polynomials to date. The challenge with using polynomials is they severely impact model accuracy [18, 29, 12]. Previous work could not successfully train DNNs with more than $11$ layers due to exploding gradients [12]. Thus, our second contribution is tailoring the ML training process to ensure high accuracy and stable training using polynomials. Our approach utilizes a new type of regularization that focuses on keeping the input to each activation function within a small range. We achieve close to plaintext accuracy on models as deep as ResNet-110 [16] and as large as a ResNet-50 on ImageNet [10] (23 million parameters). The combination of these approaches yields a co-design with state-of-the-art inference times and the highest accuracy for polynomial models.
We compare our work with three solutions representing the state-of-the-art approaches in secure inference according to Ng and Chow [32]. We summarize our results in Figure 1. Combining the single-round MPC protocols with our activation regularization achieves significantly faster inference times than all other solutions. Specifically, our solution is faster than CryptGPU by 4×4$\times$4 ×, GForce by 5×5$\times$5 ×, and COINN by 40×40$\times$40 × on average in wide area networks. Our approach also scales to large models on ImageNet with a 110×110$\times$110 × speedup over COINN, 14×14$\times$14 × speedup over Cheetah, and a 3×3$\times$3 × speedup over CryptGPU. Furthermore, our inference accuracy remains competitive with all other solutions. CryptGPU often gives slightly higher accuracy as it can evaluate any plaintext model (albeit slower than our work). Thus, the challenge for future work is to further close the ML accuracy gap between plain and polynomial models.
@@ 2.2 Neural Network Inference
We consider DNN classifiers with domain $\mathcal{X}\subseteq\mathbb{R}^{d}$ and range $\mathcal{Y}\subseteq\mathbb{R}^{c}$. DNN classifiers consist of a sequence of layers, each performing either a linear or a non-linear operation. The ResNet [16] architecture we consider is composed of (i) convolutional, (ii) fully connected, (iii) pooling, (iv) batch normalization, and (v) ReLU layers. All layers are linear, except for $\text{ReLU}(x)=\text{max}(x,0)$ and max pooling (that can be replaced with average pooling). To classify an input $x$, a classifier $h$, passes the input sequentially through each layer. Upon reaching the last layer, the prediction is obtained by taking $\text{arg max}_{i\in\{1..c\}}h(x)_{i}$, where we call $h$ the logit function for a classifier $h:\mathcal{X}\rightarrow\mathcal{Y}$. Our work is tailored to securely perform the inference phase. Specifically, for an encrypted input $x$ we compute the encrypted output $h(x)$ of the logit function. However, to achieve improvements during inference, our work requires modification to the training phase that generates $h$.
@@ 3.1 Problem Setup
We follow the same threat model as prior work for two-party secure inference [18, 22, 29, 31]. Specifically, we follow the two-party client-server model where the server has a machine learning model (a DNN) they have trained (with our technique), and the client holds private data upon which they would like to make an inference. Like Delphi [29], our work assumes the server uses a modified training procedure. The server’s input to the protocol is the weights of their trained model, which they do not want to leak to the client (due to intellectual property or protecting their MLaaS business [36]). The client has a private input (typically an image) they would like to classify using the model but do not want to leak this input or the prediction to the server. That is, the MPC function can be written as $f(\text{image},\text{model})=(label,\emptyset)$. Following previous work [18, 22, 29, 31], we consider the semi-honest model [15, §7.2.2], where adversaries do not deviate from the protocol but may gather information to infer private information. Also, in line with previous work, we assume the model architecture is known to both parties. This includes the dimensions and type of each layer, parameters such as field size used for inference, and the mean and standard deviation of the training set [22].
Our goal is to reduce the inference time as much as possible. We are willing to incur a small degradation in accuracy to achieve practical runtimes. For example, in the streaming setting, applications like spam detection are runtime-critical (a small accuracy trade-off can be tolerated to make it feasible). Since all the protocols we compare with contain pre-computation, we focus on the online phase for fair comparison. Furthermore, the online time determines the latency, which we focus on reducing in this work. We build from CrypTen [22], which does not implement pre-computation and rather assumes a third-party dealer for Beaver triplets. However, in practice, the server and client could generate the Beaver triplets in a pre-computation phase using off-the-shelf protocols [21, 33].
@@ 3.2 Privacy During Model Training
We focus only on the inference phase of machine learning. However, the privacy of the training process and training data is an orthogonal but essential problem. We recommend that the data owner take appropriate steps to protect the privacy of the model, such as training using differential privacy [2] or rounding the output of the inference. Furthermore, during training, care should be taken to protect against threats such as model stealing, which can be launched using only the inference result [19]. To summarize, the model owner learns nothing other than the fact a query was made. We ensure only the inference is revealed to the client; however, ML attacks that only require black box query access [19] must be defended against during the training process. Since we focus the effects of MPC on the runtime and accuracy of ML inference, we did not conflate this comparison with additional privacy preserving training goals. Any privacy-preserving technique would add a similar overhead (e.g., reducing the accuracy) to all approaches we evaluate.
@@ 3.3 Motivating the Co-Design of Activation Functions
Motivated by this bottleneck, several works have focused on either reducing the number of ReLUs or replacing ReLUs altogether [13, 12, 20, 24, 29, 30]. One approach is to approximate each ReLU with a high degree polynomial [12, 24]. The advantage of using polynomials is that polynomials can be computed using arithmetic shares, thus removing the need for expensive conversions and improving the total inference time. A significant challenge with polynomials is maintaining model accuracy [12, 18, 13, 20, 29].
@@ 4.2 ESPN: Exponentiating Secret Shared Values using Pascal’s triangle
We observe that, for each $i$ in the sum, party $A$ can compute $\mathbf{a}_{i}=x_{A}^{k-i}$ without needing to communicate with party $B$ (Alg. 1 line 4). Similarly, party $B$ can compute $x_{B}^{i}$ without communicating with party $A$ (Alg. 1 line 5). Finally, ($k\choose i$ can be computed by any party (or pre-computed ahead of time). For simplicity, we assign the computation of ($k\choose i$ to party $B$. Thus party $B$ computes $\mathbf{b}_{i}={k\choose i}x_{B}^{i}$.
Once each party has computed their respective vectors, we multiply $\mathbf{a}_{i}\cdot\mathbf{b}_{i}$ for each $i$ in parallel (Alg. 1 line 6). We carry out this multiplication using standard MPC protocols in one round. To use these multiplication protocols, each party must have a share of the input. We use a trivial additive secret sharing, where the other party inputs zero as their share to the protocol (Alg. 1 line 2). Finally, after the multiplication, the sum of the binomial theorem can be efficiently computed with no communication (Alg. 1 line 7).
@@
2:     $\mathbf{\mathbf{a}}=[[\mathbf{0}^{k}]]$, $\mathbf{\mathbf{b}}=[[\mathbf{0}^{k}]]$triangleright Initialize shares
3:     for $i=0:k$ do
4:         Party $A$ computes: $\mathbf{a}_{i}=\mathbf{a}_{i}+x_{A}^{k-i}$
5:         Party $B$ computes: $\mathbf{b}_{i}=\mathbf{b}_{i}+{k\choose i}x_{B}^{i}$
6:     $\mathbf{p}=BeaverMultiply(\mathbf{\mathbf{a}},\mathbf{\mathbf{b}})$$\triangleright$ Parallel Multiplication
7:     $s=\sum_{i=0}^{k}\mathbf{p_{i}}$triangleright $s=\mathbf{\mathbf{a}}\cdot\mathbf{\mathbf{b}}$
@@
4.3 Polynomial Evaluation with ESPN
Floating Point Considerations.
Our initial description of ESPNconsiders the integer domain for simplicity. Extending to floating point values is straightforward but requires rescaling (a standard practice in fixed-point arithmetic). We use CrypTen’s two-party, local truncation protocol to ensure we do not incur additional rounds. However, there is a negligible chance of an incorrect result from this truncation protocol due to wrap-around in the ring. Specifically, the probability of an incorrect result when truncating $x$ is $\frac{|x|}{2^{L}}$ where $2^{L}$ is the size of the ring [22, Appendix C.1.1]. This implies that $x$ must be small compared to the ring for this fast truncation protocol to be correct.
Polynomial Evaluation Protocol.
Using our exponentiation protocol, we show how to compute high-degree polynomials efficiently in a single round. We overview the protocol in Algorithm 2. First, in parallel, we compute all needed exponents using Algorithm 1. We recall that to ensure the correctness of truncation, we must ensure all intermediate values remain small. For simplicity, we define small to be that no intermediate scale becomes larger than twice the working precision $p$. We show the complete failure probability calculations of Algorithm 2 in Appendix E.
To ensure the values remain small after exponentiation, we create multiple scaled-down copies of the input $x$, proportional to each exponent we need to calculate. To compute $x^{i}$ we first scale $x$ down by $2^{-\bar{s}}$ where $\bar{s}=\lceil(i-2)p/i\rceil)$ and $p$ is the current working precision of $x$ (line 5). $\bar{s}$ is chosen such that $2^{p}$ becomes approximately $2^{2p/i}$ after scaling and thus approximately $2^{2p}$ after exponentiation by $i$. These are approximations as not all values of $i$ divide $(i-2)p$, so there is some error from taking the ceiling. To account for the additional factor of approximately two, we do an additional rescaling after each exponentiation in line 7. This rescaling incorporates $\bar{s}$ (which includes the ceiling function) to ensure all values are scaled back to $2^{p}$. Finally, after computing all powers, we can locally multiply the result by the coefficients (public values) and sum in line 8. We give a complete proof of correctness for Algorithm 1 and 2, including the truncation operator in Appendix E.
Algorithm 2 Polynomial Evaluation Protocol for 2-party additive secret-sharing
1:procedure $Poly$($[[x]],\alpha,n$)
2:     $\mathbf{p}_{1}=x$
3:     for $i=2:n$ do $\triangleright$ In parallel
4:         Let $\bar{s}=\lceil(i-2)p/i\rceil$triangleright Scale down factor
5:         $x^{\prime}_{i}=x\cdot 2^{-\bar{s}}$$\triangleright$ Scale down before Exp
6:         $\mathbf{p}_{i}=Exp(x^{\prime}_{i},i)$$\triangleright$ From Algorithm 1
7:         $\mathbf{p}^{\prime}_{i}=\mathbf{p}_{i}*2^{-p*(i-1)+\bar{s}\cdot i}$$\triangleright$ Scale down after Exp
8:     $y=\alpha_{0}+\sum_{i=1}^{n}\mathbf{\alpha}_{i}\cdot{p}^{\prime}_{i}$triangleright Locally dot product
9:     return $y$$\triangleright$ Secret-shares of $f(x)$
Hyperparameter Restrictions.
While Algorithm 2 is designed to keep intermediate values small, multiple hyperparameters determine the protocol’s effectiveness. Using results from Appendix E (namely the max of Theorem E.2 and Theorem E.3), we get that the probability of failure for a given truncation is bounded by
$Pr[\text{Truncation Failure}]\leq\frac{2^{n(\lceil\log_{2}{\lambda}\rceil+1)+2p}}{2^{L}}.$ (3)
For our experiments, we use CrypTen, with $L=64$ [22] and a default working precision of $p=16$. Assuming default values of $n=4$ and $\lambda=5$ we get a failure probability bound of $2^{-16}$. However, this is a pessimistic upper bound since we consider the worst-case input of ±$\pm 5$. Conversely, in our experiments, we find that the distribution of inputs follows an approximately Laplace distribution as shown in Appendix C. Thus, we observe a much smaller empirical failure probability.
While our default parameters were experimentally chosen to give high classification accuracy (Section 6, shows a minor degradation over plaintext accuracy), it is unclear if these values are optimal. For example, one approach to reducing the failure probability is to decrease $p$ from $16$ to $10$-bit (which gives a failure probability of $2^{-28}$). However, the trade-off is that the intermediate $x^{\prime}_{i}$ values will be truncated severely. For instance, with $n=4$, $x^{\prime}_{i}$ will be truncated to $5$-bits. We find that this loss of precision is too significant to simulate a ReLU function accurately. Similarly, while a higher degree polynomial might better approximate a ReLU, a higher degree will negatively affect both the failure probability and the precision loss of $x^{\prime}_{i}$. Future work could conduct an extensive hyperparameter search of all parameters to find the optimal trade-off. However, as we see in Section 6 (Tables 2-5), this would at most yield a $0.5\%$ increase in encrypted classification accuracy (PyTorch vs. CrypTen accuracy), for the worst model and dataset.
Evaluating Algorithm 2.
@@ 4.4 Alternative Single Round Protocol: HoneyBadger
Using dynamic programming, the parties can then compute any power ($x^{k}r^{0}$) using only additions of previously computed terms and powers of $r$. To compute polynomials using this protocol, we simply swap the call to $Exp$ in line 6 of Algorithm 2.
The advantage of Lu et al.’s protocol is that the communication is small (only the opening of $x-r$). The primary disadvantage is that the protocol requires a modified pre-computation phase, which is as difficult to pre-compute securely as the original problem (it is exponentiation). On the contrary, our binomial protocol uses standard Beaver triplets commonly found in MPC frameworks. There are well established protocols for efficiently computing these triplets, and the parties may already have them due to the popularity of Beaver’s trick. A more minor disadvantage of Lu et al.’s solution is that, while the protocol requires very little communication, it is not locally parallelizable as each dynamic programming step depends on the previous one. In contrast, our entire protocol can be executed in parallel.
We also consider the runtime of using HoneyBadger in Figure 2. We emphasize this is a runtime-only evaluation. Without our training algorithm in Section 5, none of the polynomial-based solutions can attain usable accuracy. We find that ESPN and HoneyBadger perform similarly in practice, with HoneyBadger gaining a slight advantage in very low network delay. Due to the pre-computation trade-offs, we will evaluate both approaches for the remainder of this work.